# Decision-making with Speculative Opponent Models

### 06 Mar 2024 v2

@@ Abstract
Opponent modeling has benefited a controlled agent’s decision-making by constructing models of other agents. Existing methods commonly assume access to opponents’ observations and actions, which is infeasible when opponents’ behaviors are unobservable or hard to obtain. We propose a novel multi-agent distributional actor-critic algorithm to achieve speculative opponent modeling with purely local information (i.e., the controlled agent’s observations, actions, and rewards). Specifically, the actor maintains a speculated belief of the opponents, which we call the speculative opponent models, to predict opponents’ actions using local observations and makes decisions accordingly. Further, the distributional critic models the return distribution of the policy. It reflects the quality of the actor and thus can guide the training of the speculative opponent model that the actor relies on. Extensive experiments confirm that our method successfully models opponents’ behaviors without their data and delivers superior performance against baseline methods with a faster convergence speed.
@@ Introduction
Recently, there has been a growing effort in applying multi-agent reinforcement learning (MARL) to address the complex learning tasks in multi-agent systems, including cooperation, competition, and a mix of both (Hernandez-Leal, Kartal, and Taylor 2019; Zhang, Yang, and Başar 2021). To help the controlled agent build knowledge of other agents, most existing works normally consider the opponents as a part of environment and naively feed the observations and actions of other agents to the controlled agent during training (Lowe et al. 2017; Foerster et al. 2018b; Rashid et al. 2018; Wei et al. 2018). However, other people argue that the controlled agent should be endowed with abilities to reason about other agents’ unknown goals and behaviors, thus coming up with opponent modeling (Albrecht and Stone 2018). Although there is already a considerable amount of research devoted to this opponent modeling problem, existing works commonly assume free access to opponents’ information (He et al. 2016; Foerster et al. 2018a; Raileanu et al. 2018; Tian et al. 2019; Papoudakis, Christianos, and Albrecht 2021). These canonical methods use the opponents’ observations and actions as the ground truth during the training of opponent models. However, the access to opponents’ real observations and actions may not be available or cheap in many cases. For instance, even in a simulator that the user has complete control, the opponent’s policy can come from third-party resources and thus, its observation format is unknown. When it comes to realistic training environments, even we have the full knowledge of opponents’ configuration, collecting the opponents’ running data induces high costs as the number of agents and the task complexity increase. In these cases, the agent must model its opponents with only locally available information (i.e., its own observations, actions, and rewards) if it wants to benefit from opponent modeling, which seems counter-intuitive. Thus, a natural question that comes to mind is: can we portray the opponents’ behaviors with purely local information?
This study proposes a speculation-based adversary modeling of the opponents, which we call the speculative opponent models, to predict opponents’ actions using local observations. In our algorithm, the agent maintains one speculative opponent model for each opponent, which receives the local information to predict the unknown opponent policy. At each time step, these opponent models take the agent’s observation as input and output the action probability distributions that represent the estimated opponents’ behavior. Based on these distributions, the agent samples multiple opponent joint actions, each of which is fed into the actor to output an action probability distribution for the agent. The final action probability distribution of the agent is the sum of the actor’s outputs weighted by the probability of corresponding opponent joint action. By this design, the agent policy is deeply coupled with the opponent models. We expect that a better opponent model can improve the decision quality because it provides the actor with more accurate information about the opponents. Therefore, it is possible to use the quality of agent policy as one kind of feedback to train the opponent models. This is the key insight of our work. To gather as much feedback as possible, we propose to use the distributional critic (Bellemare, Dabney, and Munos 2017; Dabney et al. 2018b) to model the return distribution of the agent policy. Compared to an expected return, the return distribution can provide more information. We wrap the opponent model aided actor and distributional critic as our multi-agent actor-critic framework to achieve opponent modeling with local information. To summarize, we highlight the following contributions: (i) To the best of our knowledge, this is the first work to achieve the opponent modeling with purely local information for MARL. (ii) This is the first work that sheds the light on the potential of distributional critic for guiding the training of opponent models, which can encourage future researches. (iii) We formally derive the policy gradient and opponent model gradient and propose a practical algorithm to train the actor and critic effectively. (iv) We evaluate the proposed method in different multi-agent environments. The evaluation confirms that our method successfully models reliable opponents’ policy without their data and achieves better performance with a faster convergence speed than baselines.
@@
Opponent modeling. Opponent modeling is a research topic that emerged alongside the game theory (Brown 1951). With the powerful representation capabilities of recent deep learning architectures, opponent modeling has ushered in significant progress (Albrecht and Stone 2018). Type-based reasoning methods (He et al. 2016; Hong et al. 2018; Raileanu et al. 2018; Albrecht and Stone 2017) assume that the opponent has one of several known types and update the belief using their observations obtained during training. Recursive reasoning methods (Rabinowitz et al. 2018; Yang et al. 2019; Zheng et al. 2018; Wen et al. 2019; Zintgraf et al. 2021) model the beliefs about the mental states of other agents via deep neural networks. Other approaches (Tian et al. 2019; Liu et al. 2019) regularize opponent modelling with novel objectives. However, the aforementioned works commonly assume access to the opponents’ observations and actions during both training and execution. Recent researches (Papoudakis, Christianos, and Albrecht 2020, 2021) argue that the access to opponents’ observations and actions during execution is often infeasible (e.g., in large-scale applications). They propose to learn an opponent model that only uses agents’ local information and thus, eliminate the need of opponents’ information during execution. However, their approaches still require the opponents’ true information during training. Thus, how to model an opponent’s policy when its information is unavailable during training is still an open problem. Our work presents the first attempt to solve this challenging problem.
Distributional reinforcement learning. Distributional reinforcement learning (RL) aims to model the return distributions rather than the expected return, and use these distributions to evaluate and optimize a policy (Bellemare, Dabney, and Munos 2017; Dabney et al. 2018a, b). Many studies have shown that distribution RL can achieve better performance than the state-of-art methods from the classical RL (Barth-Maron et al. 2018; Tessler, Tennenholtz, and Mannor 2019; Singh, Lee, and Chen 2020; Yue, Wang, and Zhou 2020). Motivated by the successes in single-agent scenarios, several studies (Lyu and Amato 2020; Hu et al. 2020) have extended the distributional RL method to multi-agent settings. They have used the distributional RL to alleviate the instability issue resulting from the exploration of other learning agents during training. Our work uses distributional RL to evaluate the quality of the agent policy and thus, guide the training of the opponent model that the policy relies on. Together with previous works, we show the great potential of distributional RL and can attract more research efforts to this area.
@@
Partially observable Markov games. A partially observable Markov game (POMG) (Lowe et al. 2017) of $n$ agents is formulated as a tuple $\mathcal{M}=\langle\mathcal{S},\mathbb{O},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ is a set of states describing the possible configuration of all agents and the external environment. Also, each agent $i$ has its own observation space $\mathbb{O}_{i}\in\mathbb{O}$. Due to the partial observability, in every state $s\in\mathcal{S}$, each agent $i$ gets a correlated observation $o_{i}$ based on its observation function $\mathcal{O}_{i}:\mathcal{S}\rightarrow\mathbb{O}_{i}$ where $\mathcal{O}_{i}\in\mathcal{O}$. The agent $i$ selects an action $a_{i}\in A_{i}$ from its own action space $A_{i}\in\mathcal{A}$ at each time step, giving rise to a joint action $[a_{1},\cdots,a_{n}]\in A_{1}\times A_{2}\times\cdots\times A_{n}$. The joint action then produces the next state by following the state transition function $\mathcal{T}:\mathcal{S}\times A_{i}\times\cdots\times A_{n}\rightarrow\mathcal{S}$. =$\mathcal{R}=\{r_{i}\}$ is the set of reward functions. After each transition, agent $i$ receives a new observation and obtains a scalar reward as a function of the state and its action $r_{i}:\mathcal{S}\times A_{i}\rightarrow\mathbb{R}$. The initial state $s\in\mathcal{S}$ is determined by some prior distribution $p:\mathcal{S}\rightarrow[0,1]$. Each agent $i$ aims to maximize its own total expected return $R_{i}=\mathbb{E}_{r^{t}\sim r_{i}(s_{t},a_{i}^{t}),(s_{t},a_{i}^{t})\sim\tau}\sum_{t=0}^{T}\gamma^{t}r^{t}$, where $\gamma$ is the discount factor, $r^{t}$ is its sampled reward at time step $t$, $\tau$ is the trajectory distribution induced by the joint policy of all agents, and $T$ is the time horizon. Without loss of generality, we assume that the $n$ agent can be divided into $|M|\leq n$ teams, and each team has $q$ agents with $1\leq q\leq n$. We consider the other teams as opponent agents which are controlled by a set of fixed policies. Note that a single agent can also form a team. We assume agents from the same team fully cooperate and thus share the same reward function.
@@
In the multi-agent scenarios, the actor-critic algorithm is applied to learn optimal policy for each agent. To stabilize the training, the common approach is to follow the centralized training and decentralized execution (CTDE) setting (Hernandez-Leal, Kartal, and Taylor 2019), where it commonly assumes that the agent $i$’s action-value function $Q_{i}^{\pi_{i}}(o_{1},\dots,o_{n},a_{1},\dots,a_{n})$ takes as input the joint observations and joint actions.
@@
where $s_{t}\sim P(\cdot|s_{t-1},a_{t-1}),a_{t}\sim\pi(\cdot|s_{t}),s_{0}=s,a_{0}=a$. $Z^{\pi}(s,a)$ is the return distribution, covering all sources of intrinsic randomness including reward function, state transition, stochastic policy sampling, and systematic stochasticity. In the distributional RL, we directly model the random return $Z^{\pi}(s,a)$ instead of its expectation. The distributional Bellman equation can be defined as,
@@
Our training follows the CTDE setting, i.e., we have the access to the observations and actions of the team we control. Note that, however, we do not know the observations and actions of the opponent team during the training. Figure 1 depicts the overall framework of the proposed distributional opponent model aided actor-critic (DOMAC) algorithm.
Figure 1: The DOMAC framework.
In DOMAC, each controlled agent $i$ has two components, the opponent model aided actor (OMA) (Figure 1(a)), and a centralized distributional critic (CDC) (Figure 1(c)). Suppose the agent $i$ has total $p$ opponents in the game. It models them by $p$ speculative opponent models, each corresponding to one of the opponents. For example, in Figure 1(a), $\mu_{\psi_{ik}}$ is the speculative opponent model for the opponent $k$ ($1\leq k\leq p$). The local information refers to the observation, action and reward of agent $i$. Once the agent $i$ receives a local observation, $o^{t}_{i}$ at time step $t$, each of its opponent models receive as input $o^{t}_{i}$, in addition to their ID. After processing, $\mu_{\psi_{ik}}$ outputs a distribution over the corresponding opponent $k$’s action space. From this distribution, the agent $i$ can sample the opponent $k$’s predicted action $\hat{a}^{t}_{ik}$. Then the agent policy $\pi_{\theta_{i}}$ takes the joint predicted action $\hat{a}^{t}_{i}=[\hat{a}^{t}_{i1},\dots,\hat{a}^{t}_{ip}]$ together with the observation $o^{t}_{i}$ as input, and outputs a distribution over agent $i$’s own action space $A_{i}$. Note that each sampled $\hat{a}^{t}_{i}$ results in one action probability distribution of the agent $i$. We conduct the sampling process for multiple times and aggregate the resulting action probability distributions into one final distribution from which the action $a^{t}_{i}$ is sampled.
@@
After all agents in the controlled team sample their actions, each controlled agent $i$’s CDC network $G_{\phi_{i}}$ takes as input the joint observation $\mathbf{o}^{t}$ and joint action $\mathbf{a}^{t}$ of its team, and computes a return distribution $Z_{i}(\mathbf{o}^{t},\mathbf{a}^{t})$. Note that each agent in the controlled team has an independent set of OMA and CDC, and the only shared are their observations and actions. This is critical for extracting the individual knowledge of each agent which may be heterogeneous for their different observations, current statuses, etc. In general, the rationales of DOMAC can be explained as the same closed optimization loop like that of the original actor-critic framework. OMA receives the local information with the beliefs of modeled opponents (i.e., predicted opponents’ actions). It improves the decision-making by adjusting opponent models and actively selecting the actions with higher evaluations according to CDC, while CDC, in turn, evaluates the improved policy. In the following, we first show that optimizing the OMA can be formulated as a policy gradient theorem. Then we formulate the optimization of CDC by using a quantile regression loss. Finally, we transform the theoretical findings into a concrete algorithm presented in Algorithm 1.

### 22 Mar 2024 v3

@@ Abstract
Opponent modelling has proven effective in enhancing the decision-making of the controlled agent by constructing models of opponent agents. However, existing methods often rely on access to the observations and actions of opponents, a requirement that is infeasible when such information is either unobservable or challenging to obtain. To address this issue, we introduce Distributional Opponent-aided Multi-agent Actor-Critic (DOMAC), the first speculative opponent modelling algorithm that relies solely on local information (i.e., the controlled agent’s observations, actions, and rewards). Specifically, the actor maintains a speculated belief about the opponents using the tailored speculative opponent models that predict the opponents’ actions using only local information. Moreover, DOMAC features distributional critic models that estimate the return distribution of the actor’s policy, yielding a more fine-grained assessment of the actor’s quality. This thus more effectively guides the training of the speculative opponent models that the actor depends upon. Furthermore, we formally derive a policy gradient theorem with the proposed opponent models. Extensive experiments under eight different challenging multi-agent benchmark tasks within the MPE, Pommerman and StarCraft Multiagent Challenge (SMAC) demonstrate that our DOMAC successfully models opponents’ behaviours and delivers superior performance against state-of-the-art methods with a faster convergence speed.
@@ Introduction
Recently, there has been a growing effort in applying multi-agent reinforcement learning (MARL) to address the complex learning tasks in multi-agent systems, including cooperation, competition, and a mix of both (Hernandez-Leal, Kartal, and Taylor 2019; Zhang, Yang, and Başar 2021; Yu et al. 2021; Papoudakis et al. 2020). The framework of centralized training with decentralized execution (CTDE) has drawn enormous attention, where the policy of each agent is trained with access to global information in a centralized way and executed given only the local observations in a decentralized manner. Empowered by CTDE, plenty of multi-agent RL (MARL) methods are proposed (Lowe et al. 2017; Foerster et al. 2018b; Rashid et al. 2018; Sunehag et al. 2017; Yu et al. 2022), including policy gradient-based and value-based ones. The policed-based methods, such as MADDPG(Lowe et al. 2017), COMA (Foerster et al. 2018b) and MAPPO (Yu et al. 2022) use the information (observations and actions) of the controlled agents to train the team policy. The value-based methods on the other hand, with VDN (Sunehag et al. 2017) and QMIX (Rashid et al. 2018) as the exemplars, construct a joint value function for the controlled agents. These works normally regard the opponents (if exist) as a part of the environment.
However, some people argue that the controlled agent should be endowed with abilities to reason about the opponents’ unknown goals and behaviours to benefit the decision-making, thus coming up with opponent modelling (Albrecht and Stone 2018). By successfully modelling the opponents, the agent can reliably estimate opponents’ behaviours as well as their goals, and adjust its policy to achieve optimal decision-making. Several recent works have proposed learning opponent models using deep learning architectures (He et al. 2016; Foerster et al. 2018a; Raileanu et al. 2018; Tian et al. 2019; Yang et al. 2018).
@@
Nonetheless, the existing works commonly assume access to the opponents’ observations and actions during both training and execution. Recent works relax the requirement of access to opponents’ observations and actions during execution (Papoudakis, Christianos, and Albrecht 2021, 2020). However, they still assume the opponents’ information as the ground truth for training the opponent models. We observe that access to the opponents’ real observations and actions during training may not be available or cheap in many cases. For instance, even in a simulator where the user has complete control, the opponent’s policy can come from third-party resources, and thus, its observation format is unknown. When it comes to realistic training environments, although we have the full knowledge of the opponents’ configuration, collecting the opponents’ running data induces high costs as the number of agents and task complexity increase. In these cases, the controlled agent can only rely on its locally available information (i.e., its own observations, actions, and rewards) to model its opponents if it wants to benefit from opponent modelling, which seems counter-intuitive. Thus, a natural but challenging question is: Can we portray the opponents’ behaviours with only the controlled agent’s local information?
To answer this question, this work aims to estimate the opponents’ behaviours with only the controlled agent’s local information. For this purpose, we design a novel opponent modelling workflow. That is, the controlled agent first speculates its opponents’ behaviours using opponent models conditioned on its local information; Then, the controlled agent makes decisions based on its speculation. Due to this design, the agent’s decisions are deeply coupled with the opponent models. We expect that a better opponent model can improve the decision quality because it provides the controlled agent with more accurate information about the opponents. Therefore, it is possible to use the return resulting from the controlled agent’s decisions as one kind of feedback, which takes the place of the true opponent information, to train the opponent models. This is the key insight of our work.
Following the above workflow, we propose the Distributional Opponent-aided Multi-agent Actor-Critic (DOMAC), an innovative algorithm tailored for speculative opponent modelling utilizing solely local information. The approach consists of two primary modules: the Opponent-aided Actor (OMA) and the Centralized Distributional Critic (CDC). Within the OMA, the controlled agent, i.e., the actor, maintains one opponent model for each opponent, which predicts the opponent’s actions based on the controlled agent’s local information. Although these inputs are not directly from the opponent strategies, they still correlate with the opponents’ behaviours. For example, the opponents could appear in the observations of the controlled agents, thus the local observations can occasionally convey the state-changing information of the opponents. Since our opponent modelling does not use the opponents’ information as the ground truth during training or execution, we call our opponent models the speculative opponent models. These speculative opponent models are integral to enhancing the actor’s decision-making capabilities by providing refined predictions for improved performance. Meanwhile, to gather as much feedback as possible for training the OMA, the CDC capitalizes on the concept of distributional critics  (Bellemare, Dabney, and Munos 2017; Dabney et al. 2018a, b; Bellemare, Dabney, and Rowland 2023; Lyu and Amato 2020; Hu et al. 2020; Sun, Lee, and Lee 2021; Huang et al. 2022) to appraise the distribution of the actor’s returns, rather than the conventional scalar estimate. This distributional perspective allows for the explicit consideration of correlations between values, thereby capturing more information concerning the return. It avails a more nuanced and informative evaluation of the actor’s policy that underpins the refinement of the speculative opponent models essential for the actor’s optimization. We highlight the following contributions:
  • To the best of our knowledge, we are the first work that studies the opponent modelling with purely local information from the controlled agent.
  • In the proposed DOMAC algorithm, we have tailored the designs of OMA and CDC. Besides, we are pioneering the use of distributional critics to guide the training of the opponent models, which can inspire future research.
  • Furthermore, we present a formal derivation, grounded in the policy gradient theorem, that guides the joint training of both our agent policy and opponent models, bolstering the theoretical foundation of our DOMAC designs.
  • Lastly, we empirically show that DOMAC significantly outperforms existing state-of-the-art methods on several challenging multi-agent tasks, including MPE(Lowe et al. 2017), Pommerman(Resnick et al. 2018), and StarCraft Multiagent Challenge (SMAC) (Samvelyan et al. 2019). The extensive experiments confirm that our method successfully learns reliable opponent models without using the opponents’ true information and achieves better task performance with a faster convergence speed than baselines.
The remainder of this paper is organized as follows. Section Related work briefly reviews the works in multi-agent RL, opponent modelling, and distributional RL. Section Preliminary introduces the preliminaries of the proposed method. Section Methodology elaborates on the detailed designs of the DOMAC. Section Experiments provides the evaluation experiments and ablations studies. Finally, Section Conclusion and Future Work concludes the paper and presents future works.
@@
Multi-agent reinforcement learning. In recent years, the developments in multi-agent reinforcement learning (MARL) have led to great progress in creating artificial agents that can efficiently cooperate to solve tasks (Hernandez-Leal, Kartal, and Taylor 2019; Zhang, Yang, and Başar 2021; Gronauer and Diepold 2022). MARL algorithms generally fall into two frameworks: centralized and decentralized learning. Centralized methods (Claus and Boutilier 1998) learn a single policy to produce the joint actions of all agents directly, while decentralized learning (Littman 1994) entails each agent independently optimizing its reward. The framework of centralized training and decentralized execution (CTDE) bridges the gap between the two aforementioned frameworks, which permits the sharing of information during training, while policies are only conditioned on the agents’ local observations enabling decentralized execution(Lowe et al. 2017). One category of CTDE algorithm is policy gradient methods, wherein each agent consists of a decentralized actor and a centralized critic, which is optimized by the shared information of the controlled agents(Lowe et al. 2017; Foerster et al. 2018b). The value decomposition method is another category, which represents the joint Q-function as a function of agents’ local Q-functions (Sunehag et al. 2017; Rashid et al. 2018; Son et al. 2019; Wang et al. 2020). However, these studies often consider opponents (if exist) as part of the environment, failing to explicitly model the influence of the opponents, which yields sub-optimal learning outcomes.
Opponent modelling. Opponent modelling is a research topic that emerged alongside the game theory (Brown 1951). With the powerful representation capabilities of recent deep learning architectures, opponent modelling has ushered in significant progress (Albrecht and Stone 2018).
@@
Type-based reasoning methods (He et al. 2016; Hong et al. 2018; Raileanu et al. 2018; Albrecht and Stone 2017) assume that the opponent has one of several known types and update the belief using their observations obtained during training. Recursive reasoning methods  (Yang et al. 2019; Wen et al. 2019b, a) model the beliefs about the mental states of other agents via deep neural networks. The opponent modelling with online reasoning refers to the online inference of the opponent’s policy through Bayesian inference and formulating a corresponding response (Zheng et al. 2018; Tian et al. 2019; DiGiovanni and Tewari 2021; Fu et al. 2022). The process of modelling opponents with dynamic strategies involves estimating the opponents’ behaviours and corresponding impacts by using the opponent’s learning (Foerster et al. 2017; Grover et al. 2018; Kim et al. 2021; Liu et al. 2019). The opponent modelling with meta-learning is designed to train against a set of known opponent policies to quickly adapt to unknown opponent policies during the testing (Zintgraf et al. 2021; Wu et al. 2022). Theory of mind-based opponent modelling reasons about opponent’s mental status and intentions, to predict and adapt to opponent behaviour by modelling their beliefs, goals, and actions (Yang et al. 2018; Rabinowitz et al. 2018). However, the aforementioned works commonly assume access to the opponents’ observations and actions during both training and execution. Recent research  (Papoudakis, Christianos, and Albrecht 2020, 2021) argues that access to opponents’ observations and actions during execution is often infeasible (e.g., in large-scale applications). They manage to learn an opponent model that only uses the local information of the agent, such as its observations, actions, and rewards during execution. Nevertheless, the efficacy of their methods is constrained by the necessity of accessing opponents’ genuine data during the training phase, a requirement that significantly restricts their applicability in situations where such information is unobservable. Thus, how to model an opponent’s policy when its information is unavailable during training is still an open research problem. Our work presents the first attempt to solve this challenge.
Distributional reinforcement learning. Distributional reinforcement learning (RL) aims to model the return distributions rather than the expected return and uses these distributions to evaluate and optimize a policy  (Bellemare, Dabney, and Munos 2017; Dabney et al. 2018a, b; Bellemare, Dabney, and Rowland 2023). Many studies have shown that distribution RL can achieve better performance than the state-of-art methods from the classical RL  (Barth-Maron et al. 2018; Tessler, Tennenholtz, and Mannor 2019; Singh, Lee, and Chen 2020; Yue, Wang, and Zhou 2020; Bai et al. 2023; Zhang et al. 2023; Wiltzer, Meger, and Bellemare 2022). The Categorical DQN (C51) (Bellemare, Dabney, and Munos 2017) employed a discrete set of $N$ fixed values to approximate the return distribution, achieving superior performance than DQN. Then, Dabney et al. introduced QR-DQN (Dabney et al. 2018b), by appealing to the principle of quantile regression (QR) (Koenker and Hallock 2001), which assigns $N$ fixed and uniform probabilities to adjustable values. Furthermore, an extension to QR-DQN was realized through implicit quantile networks (IQNs) (Dabney et al. 2018a), which learn the entire quantile function via neural networks. Interestingly, the latest research in Nature also shows that similar distributional mechanisms exist in human brains (Dabney et al. 2020), motivating a variety of hybrid methodologies to blend these distributional methods with existing RL techniques to address issues within single-agent (Barth-Maron et al. 2018; Ma et al. 2020) and multi-agent scenarios (Lyu and Amato 2020; Hu et al. 2020; Sun, Lee, and Lee 2021; Huang et al. 2022).
Our work is the first to investigate distributional RL for opponent modelling. Together with previous works, we show the great potential of distributional RL and can attract more research efforts to this area.
@@
Partially observable Markov games. A partially observable Markov game (POMG) (Lowe et al. 2017) of $n$ agents is formulated as a tuple $\mathcal{M}=\langle\mathcal{S},\mathbb{O},\mathcal{O},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ is a set of states describing the possible configuration of all agents and the external environment. Also, each agent $i$ has its own observation space $\mathbb{O}_{i}\in\mathbb{O}$. Due to the partial observability, in every state $s\in\mathcal{S}$, each agent $i$ gets a correlated observation $o_{i}$ based on its observation function $\mathcal{O}_{i}:\mathcal{S}\rightarrow\mathbb{O}_{i}$ where $\mathcal{O}_{i}\in\mathcal{O}$. The agent $i$ selects an action $a_{i}\in A_{i}$ from its own action space $A_{i}\in\mathcal{A}$ at each time step, giving rise to a joint action $[a_{1},\cdots,a_{n}]\in A_{1}\times A_{2}\times\cdots\times A_{n}$. The joint action then produces the next state by following the state transition function $\mathcal{T}:\mathcal{S}\times A_{i}\times\cdots\times A_{n}\rightarrow\mathcal{S}$. =$\mathcal{R}=\{r_{i}\}$ is the set of reward functions. After each transition, agent $i$ receives a new observation and obtains a scalar reward as a function of the state and its action $r_{i}:\mathcal{S}\times A_{i}\rightarrow\mathbb{R}$. The initial state $s\in\mathcal{S}$ is determined by some prior distribution $p:\mathcal{S}\rightarrow[0,1]$. Each agent $i$ aims to maximize its own total expected return $R_{i}=\mathbb{E}_{r^{t}\sim r_{i}(s_{t},a_{i}^{t}),(s_{t},a_{i}^{t})\sim\tau}\sum_{t=0}^{T}\gamma^{t}r^{t}$, where $\gamma$ is the discount factor, $r^{t}$ is its sampled reward at time step $t$, $\tau$ is the trajectory distribution induced by the joint policy of all agents, and $T$ is the time horizon. Without loss of generality, we assume that the $n$ agent can be divided into $|M|\leq n$ teams, and each team has $q$ agents with $1\leq q\leq n$. We consider the teams out of our control as opponent agents with static policies. Note that a single agent can also form a team. We assume agents from the same team fully cooperate and thus share the same reward function.
@@
In multi-agent scenarios, the actor-critic algorithm is applied to learn the optimal policy for each agent. To stabilize the training, it is common to adopt the centralized training and decentralized execution (CTDE) setting (Hernandez-Leal, Kartal, and Taylor 2019), where the agent $i$’s action-value function $Q_{i}^{\pi_{i}}(\mathbf{o},\mathbf{a})$ takes as input the joint observation and joint action of the agents in the same controlled team. The gradient of the policy for agent $i$ is given as:
$\nabla_{\theta_{i}}J(\theta_{i})=\mathbb{E}_{\pi_{\theta_{i}}}[\nabla_{\theta_{i}}\log\pi_{\theta_{i}}(\mathbf{a}|\mathbf{o})Q_{\phi}^{\pi}(\mathbf{o},\mathbf{a})].$ (2)
@@
where $s_{t}\sim P(\cdot|s_{t-1},a_{t-1}),a_{t}\sim\pi(\cdot|s_{t}),s_{0}=s,a_{0}=a$. $Z^{\pi}(s,a)$ is the return distribution, covering all sources of intrinsic randomness, including reward function, state transition, stochastic policy sampling, and systematic stochasticity. In the distributional RL, we directly model the random return $Z^{\pi}(s,a)$ instead of its expectation. The distributional Bellman equation can be defined as:
@@
In general, the distribution of the random variable $Z(s,a)$ is represented as $\frac{1}{K}\sum_{j=1}^{K}{\delta_{G_{\phi}^{j}(\mathbf{o},\mathbf{a})}}$ where $\delta_{G_{\phi}^{j}(\mathbf{o},\mathbf{a})}$ denotes a Dirac with respect to the support $G_{\phi}^{j}(\mathbf{o},\mathbf{a})\in\mathbb{R}$. Furthermore, each $G_{\phi}^{j}(\mathbf{o},\mathbf{a})$ can be estimated through optimizing the quantile Huber loss (Dabney et al. 2018a):
$\mathcal{L}_{QR}=\frac{1}{K^{2}}\sum_{j^{\prime}=1}^{K}\sum_{j=1}^{K}[\rho_{\omega_{j}}(\delta_{ij})],$ (5)
where $\rho_{\omega_{j}}(\mu)=|\omega_{j}-\mathbbm{1}_{\mu\leq 0}|\mathcal{L}_{\kappa}(\mu)$ and $\delta_{ij}=r+\gamma G_{j}(\mathbf{o}^{\prime},\mathbf{a}^{\prime})-G_{j}(\mathbf{o},\mathbf{a})$. $\mathbbm{1}_{\mu\leq 0}$ is an indicator function which is $1$ when $\mu\leq 0$ and 00 otherwise. $\kappa$ is a pre-fixed threshold and the Huber loss $\mathcal{L}_{\kappa}(\mu)$ is given by,
Table 1: Notations of algorithm
Notation Meaning
$p$
The numbers of opponents
$\mu_{\psi_{ik}}$
Predicted opponent model for opponent $k$ of agent $i$
$o_{i}$
Local observation of agent $i$
$\hat{a}_{ik}^{t}$
Predicted opponent actions for opponent $k$ of agent $i$
$\pi_{\theta_{i}}$
Policy network of agent $i$
$a_{i}^{t}$
Action of agent $i$
$\rho_{\theta_{i},\psi_{i}}$
Final agent policy distribution
$Z_{i}(\mathbf{o}^{t},\mathbf{a}^{t})$
Return distribution of agent $i$
$G_{\phi_{i}}$
CDC network
$\omega$
Quantiles
$\delta_{z}$
Dirac at $z\in\mathbb{R}$
$K$
Quantile number
$\alpha_{i}$
hyperparameter of exploration
@@
Our training follows the CTDE setting, i.e., we have the access to the observations and actions of the team we control. Note that, however, we do not know the observations and actions of the opponent team during the training. Figure 1 depicts the overall framework of the proposed distributional opponent model aided actor-critic (DOMAC) algorithm. We show the meaning of the notifications in Table 1.
Within the DOMAC framework, each controlled agent $i$ incorporates two primary constituents: the Opponent Model Aided Actor (OMA) and the Centralized Distributional Critic (CDC). For $p$ opponents, agent $i$ employs $p$ individualized conjectural opponent models, wherein each model is instantiated as a distinct neural network. These models process the agent’s localized observational data and sample actions at each timestamp to ascertain the presumptive action distributions of the opponents. The OMA leverages these models to infer potential joint actions, representing the agent’s conjectures concerning the opponents’ prospective moves. The agent $i$’s own action distribution is subsequently determined as a composite function of the outputs from the OMA and the observations, effectively assimilating the agent’s tactical considerations with the inferences drawn from the opponent models by performing a weighted aggregation. To enhance predictive precision, the agent samples multiple opponents’ joint actions. Intriguingly, the opponent models’ training regimen is predicated on the rewards obtained by the agent’s policy alone, eschewing reliance on the opponents’ genuine actions and underscoring outcome-oriented feedback loops. Meanwhile, the CDC’s role is to evaluate the return distribution associated with the agent’s policy, preferring a distributional perspective over singular expected value estimations. This methodological choice for estimating returns confers a more nuanced depiction, which polishes the refinement process for the agent’s strategic policy.
@@
Note that each agent in the controlled team has an independent set of OMA and CDC, and the only shared are their observations and actions. This is critical for extracting the individual knowledge of each agent, which may be heterogeneous for their different observations, current statuses, etc. To speed up the learning, we share the parameters of the agent networks across all controlled agents but a one-hot encoding of the $agent_{id}$ is concatenated onto each agent’s observations to induce heterogeneities of the inputs. In scenarios where the agents within a system are homogeneous, both functionally and behaviorally, the utilization of a shared OMA network is feasible for these homogeneous agents. Generally, the rationales of DOMAC can be explained as the same closed optimization loop as that of the original actor-critic framework. That is, OMA receives the local information with the beliefs of modelled opponents (i.e., predicted opponents’ actions). It improves decision-making quality by adjusting opponent models and actively selecting the actions with higher evaluations according to CDC, while CDC, in turn, evaluates the improved policy. In the following, we first show that optimizing the OMA can be formulated as a policy gradient theorem. Then we formulate the optimization of CDC by using a quantile regression loss. Finally, we transform the theoretical findings into a practical algorithm.
Figure 1: The DOMAC framework. In OMA, the speculative opponent model $\mu_{i}$ takes the local observation $o^{t}_{i}$ as input and outputs the prediction of the opponent behaviours. The agent policy network takes action by considering the predicted opponent’s actions. The CDC takes the joint observation $\mathbf{o}^{t}$ and action $\mathbf{a}^{t}$ of the controlled agents into the agent critic network $G_{\phi_{i}}$ and outputs the return distribution $Z_{i}(\mathbf{o}^{t},\mathbf{a}^{t})$.
Figure 2: An illustration of our DOMAC network architecture. (a) Opponent model-aided actor: Each model contains $p$ speculative opponent models, which take the local observation $o^{t}_{i}$ and opponent index $k$ as inputs to predict the opponents’ actions. Then, the action selection network $\pi_{\theta_{i}}$ takes the joint predicted action $\{\hat{a}^{t}_{i}\}$ together with the $o^{t}_{i}$ and $a^{t-1}_{i}$ as input, and outputs a distribution over the agent $i$’s own actions, which is weighed according to the probabilities of predicted opponents’ actions $\mu^{k}_{\phi_{i}}$ for $1\leq k\leq p$. (b) CTDE architecture: we have the access to the observations and actions of the team we control. (c) Centralized distributional critic: the agent critic network takes the joint observation $\mathbf{o}^{t}$ and action $\mathbf{a}^{t}$ of the controlled agents and outputs $G_{\Phi_{i}}$ as the approximation of return distribution.